OpenStudy (anonymous):

Fan/medal will be rewarded How do you find the solution of an equation that has "no real solutions" from a negative discriminant.

OpenStudy (helder_edwin):

u can use the quadratic formula, but since the discriminant is negative u can factor the sign inside the square root \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{-b\pm\sqrt{(-1)(4ac-b^2)}}{2a} \] \[ \large =\frac{-b\pm\sqrt{4ac-b^2}\sqrt{-1}}{2a} \]

OpenStudy (ikram002p):

u just but i insted of sqart -1,and solve like if it was reall. any way do have a specific prob ??

OpenStudy (helder_edwin):

now the number inside the square root is positive and u get two complex solutions

OpenStudy (anonymous):

the problem is -x\[^{2}\] +3x-6=0 @ikram002p

OpenStudy (helder_edwin):

use the quadratic formula

OpenStudy (anonymous):

I'm still confused

OpenStudy (helder_edwin):

ok u have \[ \large -x^2+3x-6=0 \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

so a=-1, b=3, and c=-6 right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

so \[ \large x=\frac{-3\pm\sqrt{3^2-4(-1)(-6)}}{2(-1)} \] ok?

OpenStudy (anonymous):

Okay

OpenStudy (helder_edwin):

so \[ \large x=\frac{-3\pm\sqrt{9-24}}{-2}=\frac{-3\pm\sqrt{-15}}{-2} \] right?

OpenStudy (anonymous):

Okay, I'm with you so far.

OpenStudy (helder_edwin):

now \[ \large x=\frac{-3\pm\sqrt{15(-1)}}{-2}=\frac{-3\pm\sqrt{15}\sqrt{-1}}{-2} =\frac{-3\pm i\sqrt{15}}{-2} \] where \(i=\sqrt{-1}\).

OpenStudy (anonymous):

Okay, still with you.

OpenStudy (helder_edwin):

that is it. 15 is not a perfect square.

OpenStudy (anonymous):

That's it?

OpenStudy (helder_edwin):

yes

OpenStudy (ikram002p):

Amanda hope u the best :) helder good job lol.

OpenStudy (helder_edwin):

did u understand? @AmandaMichelle

OpenStudy (anonymous):

Thanks @ikram002p and yes it all makes sense now thank you :) @helder_edwin

OpenStudy (helder_edwin):

u r welcome